OpenStudy (anonymous):

Solve for s. -2s – 19 = 6s + 13

OpenStudy (anonymous):

@MollyNicole2015

OpenStudy (anonymous):

\[-2s-6s=13+19\]\[-8s=32\]

OpenStudy (mollynicole2015):

Simplifying -2s + -19 = 6s + 13 Reorder the terms: -19 + -2s = 6s + 13 Reorder the terms: -19 + -2s = 13 + 6s Solving -19 + -2s = 13 + 6s Solving for variable 's'. Move all terms containing s to the left, all other terms to the right. Add '-6s' to each side of the equation. -19 + -2s + -6s = 13 + 6s + -6s Combine like terms: -2s + -6s = -8s -19 + -8s = 13 + 6s + -6s Combine like terms: 6s + -6s = 0 -19 + -8s = 13 + 0 -19 + -8s = 13 Add '19' to each side of the equation. -19 + 19 + -8s = 13 + 19 Combine like terms: -19 + 19 = 0 0 + -8s = 13 + 19 -8s = 13 + 19 Combine like terms: 13 + 19 = 32 -8s = 32 Divide each side by '-8'. s = -4 Simplifying s = -4

OpenStudy (mollynicole2015):

Now what do you do to verify it's right?

OpenStudy (anonymous):

Put the answer back into the equation.

OpenStudy (confxsedaf):

@MollyNicole2015 is correct :)

OpenStudy (mollynicole2015):

Right! Do it.

OpenStudy (mollynicole2015):

:) Thanks

OpenStudy (anonymous):

-2 x -4 – 19 = 6 x -4 + 13 -2 x -4 - 19 = -11 6 x -4 + 13 = -11 -11 = -11

OpenStudy (mollynicole2015):

Good! So that means what?

OpenStudy (anonymous):

Our equation is correct.

OpenStudy (mollynicole2015):

Very good!

OpenStudy (anonymous):

I'm going to medal you because you're so good at helping. :)

OpenStudy (mollynicole2015):

Thank you. <3 And you deserve a medal for your hard work!